### v1

# GraphCompNet: A Position-Aware Model for Predicting and Compensating Shape Deviations in 3D Printing

Abstract
Shape deviation modeling and compensation in additive manufacturing (AM) are pivotal for achieving high geometric accuracy and enabling industrial-scale production. While traditional analytical and statistical methods laid the foundation, recent advancements in machine learning (ML) have improved prediction and compensation precision. However, critical challenges persist, including generalizability across complex geometries and adaptability to position-dependent variations in batch production. Traditional methods of controlling geometric deviations often rely on complex parameterized models and repetitive metrology, which can be time-consuming yet not applicable for batch production. In this paper, we present a novel approach to address this challenge of ensuring geometric precision and accuracy in position-dependent powder bed fusion (PBF) production. The proposed GraphCompNet presents a novel computational framework integrating graph-based neural networks with a generative adversarial network (GAN)-inspired training paradigm. The framework leverages point cloud representations and dynamic graph convolutional neural networks (DGCNNs) to model intricate geometries while incorporating position-specific thermal and mechanical variations. A two-stage adversarial training process iteratively refines compensated designs using a compensator-predictor architecture, enabling real-time feedback and optimization. Experimental validation across various shapes and positions demonstrates the framework’s ability to predict deviations in freeform geometries and adapt to position-dependent batch production conditions, significantly improving compensation accuracy (35% to 65%) across the entire printing space, addressing position-dependent variabilities within the print chamber. The proposed method advances the development of a Digital Twin for AM, offering scalable, real-time monitoring and compensation capabilities. This work bridges critical gaps in AM process control, paving the way for high-precision, automated, and industrial-scale design and manufacturing systems.
Index Terms:
Additive Manufacturing, Powder-bed Fusion, Shape Deviation, Shape Compensation, Graph Neural Network, Quality control, Digital Twin
publicationid: pubid: $0$–$00$ 2021 IEEE
Note to Practitioners—This paper introduces a framework for predicting and compensating position-dependent shape deviations in 3D printing. The proposed approach effectively models complex and arbitrary 3D geometries while addressing variabilities across different printer positions, making AM more suitable for industrial batch production. This advancement demonstrates the potential to integrate Digital Twin technology into AM processes, enabling closed-loop design optimization and enhancing both precision and scalability in large-scale, practical applications. The code for this framework is available on the NVIDIA Modulus platform: https://github.com/NVIDIA/modulus.
Figure 1: Configuration of the Molded Fiber dataset within the printing chamber, illustrating part placement, including stacking in the y-orientation and rotation in the x-orientation. The top row displays the deviations of two identical parts placed at different positions within the same print bucket, with a heatmap indicating the scale of deviation. These parts display distinct deviation patterns across their geometries, as highlighted by the varying color map distributions. The bottom row shows the decreasing deviations after compensation using the proposed compensation framework in both parts.
I Introduction
Powder Bed Fusion (PBF) is an advanced additive manufacturing (AM) technique that fuses powdered materials—typically metals, plastics, or ceramics—using a laser or electron beam. By precisely controlling energy deposition at the voxel level, PBF creates highly complex and detailed structures with fine resolution (80-250 microns). It is widely used in industries like aerospace and biomedicine for high-quality output and the ability to produce multiple parts in a single build. Among PBF methods, Multi Jet Fusion (MJF) [1] improves production efficiency by injecting a rapid agent at the pixel level and projecting thermal energy across the build area, enabling faster print speeds without compromising accuracy or quality. A critical challenge in AM in general, however, is achieving consistent geometric accuracy.
For instance, as shown in Fig. 1, a molded fiber part placed at the center of the print chamber (blue in Fig. 1) demonstrates a maximum deviation of 6 mm from the nominal design, with a standard deviation of 1 mm. In contrast, a part placed at the bottom of the chamber (pink in Fig. 1) shows a maximum deviation of 7 mm and a standard deviation of 1.1 mm. These parts display distinct deviation patterns across their geometries, as highlighted by the varying color map distributions. Achieving high precision in AM is challenging due to the complex interactions of multiple factors, including material properties, part orientation, printer settings, calibration, layer-by-layer deposition, energy distribution, temperature gradients, and residual stresses, among others. Early research focused on analytical and statistical models to predict and mitigate these deviations, but their reliance on simplified assumptions limited applicability to real-world geometries. Recent advances in computational methods, particularly machine learning (ML), have introduced data-driven models that enhance prediction accuracy and scalability. Despite progress, critical challenges hinder the transition of these models to industrial-scale AM.
Existing deviation prediction models face challenges with generalizing across complex geometries. FEA-based models [2] are computationally expensive for real-time use, while data-driven frameworks for 3D deviation prediction are often validated only for simple shapes (e.g., domes). Furthermore, compensation models frequently overlook position-dependent variations inherent in batch production (referred to as ”position-aware” below), such as the significant deviations caused by spatial factors like thermal gradients. Additionally, the lack of experimental validation in prior research limits the practical applicability of these models in industrial settings.
Since batch production is essential for industrial applications, overcoming these challenges is crucial for developing more efficient, automated systems like a Digital Twin, which can improve precision, control, and scalability in AM. A breakthrough in this area would significantly enhance both design accuracy and mass production reliability, driving further innovation in the field. The remainder of the paper is structured as follows: Section II reviews deviation prediction and compensation models, as well as the research challenges, Section III details the proposed methodology, Section IV presents experimental results, and Section V concludes with future directions.
II State-of-the-art and research challenges
Shape deviation modeling and compensation in AM have been critical for improving geometric accuracy and quality control. Initially, research focused on analytical and statistical models, recent advancements have incorporated computational techniques such as machine learning approaches. This section highlights three key areas: (1) review of Deviation Prediction Modeling, (2) review of Deviation Compensation Modeling, and (3) Research Challenges and Our Proposed Methodology, emphasizing the transition to data-driven strategies and the hurdles in achieving industrial-scale precision in AM mass production.
II-A Deviation Prediction Modeling
Early research on shape deviation often employed statistical and analytical models for parameterization, both in 2Dand 3D [3, 4, 5]. Schmutzler et al. [6] innovated by framing shape deviation prediction as a non-rigid shape registration problem. This approach, utilizing B-spline-based registration, predicted deviations and compensated for CAD parts by shifting CAD vertices in the opposite direction of the learned deviation [7]. Huang et al. [8] developed statistical models using Polar and Spherical Coordinate Systems (PCS, SCS). These models address predicting in-plane and out-of-plane deviations, though their reliance on experimental setups and calibration limited scalability and adaptability. Additionally, finite element analysis (FEA) was applied to simulate part deformation based on real process instructions .
More recently, machine learning (ML) techniques have been incorporated into deviation prediction to improve accuracy. Ferreira et al. [9] used Bayesian neural networks to model and compensate for deviations. Decker et al. [10] developed a random forest model for 3D freeform shapes, enabling rapid, minimal human intervention. Convolutional neural networks (CNNs) [11] and Shape Deviation Generators [12] have been applied to detect complex deviation patterns, while Wang et al. [13] extended ML to both smooth and non-smooth shapes. Li et al. [14] proposed in-situ deviation monitoring during printing.
Despite advances, several challenges remain. Analytical and statistical models are computationally expensive, with analyses on a single part often taking hours. While ML methods show promise, they are often limited to simpler geometries (such as simple freeform shapes based on cylinders and polyhedrons), leaving a gap for more intricate 3D geometries. This highlights the need for more efficient and generalized models capable of addressing both computational constraints and the complexity of real-world AM geometries.
II-B Deviation Compensation Modeling
[…]
To address these challenges, advanced algorithms have been introduced. Bayesian models [9] use uncertainty quantification to correct for geometric errors in stereolithography printing, improving robustness across different printer setups. Decker et al. [10] applied random forest models to predict deviations in fused deposition modeling (FDM) with a small training dataset, demonstrating over 44% shape deviation. Though domain knowledge based on the geometry is essential to define the predictor variables for the random forest model. Hong et al. [19] used artificial neural networks (ANN) to enhance dimensional accuracy in truss lattice structures, outperforming traditional methods in selective laser melting. The method highlighted the potential of using a data-driven approach to generate flexible, free-form geometries, providing enhanced compensation. Despite recent advancements, existing compensation techniques generally rely on shape deviation models for isolated individual shapes. The substantial effect of thermal processes on shape deviations requires modeling parts in different positions within the printer, which demands considerable resources for creating position-specific deviation models, even for a single geometry.
II-C Research Challenges and Proposed Methodology
Despite significant advancements in shape deviation modeling and compensation, key challenges persist, particularly regarding scalability, data dependency, and generalization. These challenges emphasize the need for novel approaches that address these gaps, ultimately enhancing the efficiency and accuracy of AM. This research proposes a computational framework to predict and compensate for print geometry, incorporating position-specific variables such as energy differences in the print chamber of the same print geometry. Limited research has focused on addressing position-specific compensation in AM, making this contribution vital. The primary challenges in this context are:
Figure 2: The configuration of the bar dataset within the printing chamber (a) illustrates the distribution of identical part geometries placed at different positions for batch production, each labeled with a part ID. The selected training and validation parts for the case studies are highlighted with circles, as shown in (b), a cross-section of the build bed layout.
  • Generalizability Across Complex Printing Geometries: Exisiting works demonstrate modeling, predicting and compensating for 2D shape deviations [20, 21]. However, modeling and control 3D geometries taking into account the layer-wise interactions presents significantly greater challenges. Huang et al. [12, 13] introduced a data-analytical framework for 3D deviation, but its validation has been limited to simple shapes such as domes and stacked rectangles. Approximating non-smooth 3D freeform shapes using a combination of patches remains constrained by resolution limits and requires additional steps to subtract polyhedral shapes. Additionally, Finite Element Analysis (FEA)-based models are computationally expensive and impractical for complex industrial geometries and batch processes [22]. More recent methods using Convolutional Neural Networks (CNNs) [23, 11] for error prediction. Shen et al. [24] employed an Encoder-Decoder architecture to predict and compensate binary image data. However, those methods are limited by their resolution and lack of detail, due to preprocessing in fixed-size images or binary format.
  • Adaptability to Position-Dependent Batch Production: Traditional and machine learning-based compensation models often fail to consider the dynamic thermal and mechanical factors that influence AM, leading to suboptimal results under varying conditions [25]. Thermal-induced deformations significantly impact part distortion. Research by Chen et al. [26] shows how thermal variations across the print bed result in significant differences in part quality. For example, Figure 2 illustrates a dataset consisting of 140 identical bars placed at different positions within the printing bed. The shape deviation patterns exhibit significant variation, with parts positioned at the edges of the print chamber, such as Part ID 140 and Part ID 31 in Fig. 2 (b), experiencing greater deviations compared to those located at the center, such as Part ID 72 and Part ID 24. Hartmann et al. [27] highlighted the importance of batch production in industrial settings, demonstrating a data-driven compensation approach applied to stacked fin-shaped parts. However, their method is constrained by the need for iterative redesigns and was validated using only a single geometry.
    This research identifies this gap in experimental validation, mainly due to high costs and resource-intensive processes. The importance of position-specific deviation modeling within the printer is often overlooked, making it a significant challenge for automation in industrial batch production. The lack of position awareness in current models is a critical barrier to implementing a Digital Twin for AM, which is necessary for improving precision and control throughout the industrial production process.
The proposed method
To address the two key challenges outlined above, we introduce a computational framework leverages graph data structure-based neural network architecture to learn position-specific information and model shape deviations from print scan data and corresponding CAD files. The proposed Compensator-Predictor learning framework (GraphCompNet), inspired by Generative Adversarial Networks (GANs), enables the generation of compensated designs directly for any part position within the print chamber.
Figure 3: The schematic diagram illustrates the proposed architecture for shape deviation prediction and compensation, inspired by the generative adversarial network framework. Initially, the DL compensation engine generates a compensation plan (step 1), which is subsequently evaluated by the DL prediction engine through shape deviation prediction (step 2). Finally, the process iterates as needed, incorporating data-driven feedback (step 3).
  1. 1.
    To address the first challenge, to enable shape deviation prediction and compensation for arbitrary, complex geometries, we introduce two innovations:
    1. (a)
      Point Cloud Data Representation: A major challenge in modeling shape deviation and compensation for complex geometries stems from the limitations of structured data formats like images or fixed 3D structures (e.g., voxels and meshes). These formats struggle with resolution constraints for inputs of varying sizes and computational inefficiencies due to their discretized nature. To overcome these issues, we adopt point cloud data, which allows for more efficient handling of irregular and sparse data typically encountered in industrial 3D printing. Point clouds enable the model to effectively manage variations such as object orientation, scale, and differing local densities found in real-world geometries. Additionally, point clouds naturally capture spatial relationships between data points, making them well-suited for topological pattern analysis—a key element in accurately predicting shape deviations in complex geometries.
    2. (b)
      DGCNN (Dynamic Graph Convolutional Neural Network): To further enhance the model’s ability to handle the complexities of point cloud data, we incorporate DGCNN as the backbone architecture. DGCNN dynamically constructs graph structures by connecting points based on local geometric relationships, allowing it to effectively capture the varying densities and topologies inherent in complex shapes. The EdgeConv operation within DGCNN adapts to these local variations, facilitating the aggregation of point features and improving the model’s capacity to identify fine-grained spatial patterns. This enables the model to achieve higher performance in tasks that require precise understanding of local geometry and complex spatial relationships, crucial for accurate shape prediction and deviation analysis in diverse 3D structures.
  2. 2.
    To address the second challenge, we incorporate a two-step approach to adapt to position-dependent factors:
    1. (a)
      Incorporation of Part Position Information: First, we incorporate part position information into the input geometries. For parts translated along the x and y axes, we adjust the corresponding point cloud coordinates to account for the scale. For parts that are rotated, we apply a rotation matrix to the point cloud coordinates. As a result, the input data integrates both the part geometry and placement information. This enables the model to learn position-specific variables during training, allowing it to adapt to variations in print conditions and ensuring accurate predictions, even in the presence of environmental or process-related differences.
    2. (b)
      Adversarial Training Framework:Inspired by Generative Adversarial Networks (GANs), we propose a novel approach where the generation of compensation plans and the evaluation of new geometries occur simultaneously during training. This dual-process framework allows the model to effectively generate and assess results in real-time. Specifically, we introduce two networks: one learns to generate geometries that closely resemble the optimal compensated design, while the counterpart network identifies designs that will deform into the ideal outcome. This adversarial setup encourages continuous refinement of the generated geometries.
Based on the findings above, this paper addresses Challenge 1 by leveraging point clouds to efficiently represent the varying densities and complex topologies inherent in industrial printing geometries, with adaptable resolution requirements. Additionally, backbone network architectures, such as DGCNN, are utilized to process point cloud data, facilitating the efficient learning of both local and global geometric relationships from the input geometries. Furthermore, this paper addresses Challenge 2 by directly learning position-based variabilities from the preprocessed input data. The GraphCompNet framework continuously proposes new geometries and assesses the proposed compensated geometries to meet the print deformation conditions. This approach overcomes the inherent limitation of machine learning models, which typically require large datasets for training, and reduces the need for extensive experimental validations. By leveraging position-aware inputs and adversarial training, the proposed framework enables closed-loop training and inference for batch production. This not only improves the efficiency and accuracy of the compensation process but also lays the foundation for implementing a Digital Twin for AM production, offering enhanced precision and control over the entire production process.
III Methods
Review of Geometric Deep Learning. Convolutional neural networks (CNNs) have spearheaded significant advancements in computer vision and natural language processing However, the inherent non-Euclidean nature of 3D designs presents a challenge for directly implementing standard CNNs on CAD parts. Graph neural networks (GNNs) have garnered increasing attention for geometric applications, offering a framework that accommodates non-Euclidean data structures. In GNNs, graph structures $G=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ and $\mathcal{E}$ represent the sets of vertices and edges, are leveraged to define convolution operations tailored to the given data.
Graph neural networks can be broadly categorized into two families: spectral and spatial. The spectral family of GNNs is based on the concept of generalized Fourier transforms, utilizing eigen-decomposition of the graph Laplacian [28] for convolution operations. However, computing the graph Laplacian can be computationally expensive. To address this challenge, alternative approaches have been proposed, including approximating spectral filters using polynomial filters [29]. Spatial approaches define convolution operations $h_{\theta}$ with regard to edge information between vertices $e_{ij}\in\mathcal{E}$, offering a computationally efficient alternative [30]:
$x_{i}^{(l)}=\Sigma_{j\in\mathcal{E}_{i}}h_{\theta}^{(l-1)}(x_{i}^{(l-1)},x_{j}^{(l-1)})$ (1)
where $l$ denotes $l$-th convolutions, $\theta$ is a set of learnable parameters. As a representation, Edge Convolution (EdgeConv) concatenates and aggregates features from $x_{i}^{(l)}$ and $x_{i}^{l}-x_{j}^{(l)}$, representing edge information with relative distance between neighboring vertices [31]. MoNet [30] projects neighboring points onto a pseudo-coordinate plane, then define convolution operations on projected points on the plane. Overall, the evolution of graph neural networks presents a promising path for tackling geometric applications, offering flexibility and scalability in handling non-Euclidean data structures inherent in 3D designs [32].
Review of Generative adversarial network (GAN). Generative adversarial networks (GANs) have become a potent method for modeling intricate data distributions from low-dimensional latent spaces. Recent advancements in generative deep learning have resulted in a proliferation of innovative applications across various domains, including fashion, graphic design, art, architecture, and urban planning. The GAN framework consists of two neural network components: the generator and the discriminator.
The generator maps a fixed noise distribution $\mathbb{P}_{z}$ to the generated data distribution $\mathbb{P}_{G}$, while the discriminator distinguishes samples from the distributions of generated data $\mathbb{P}_{G}$ and real data $\mathbb{P}_{r}$. GANs are trained iteratively in a minimax game, aiming to optimize the following objective function:
$\min_{G}\max_{D}\mathbb{E}_{X\sim\mathbb{P}_{r}}[\log D(X)]+\mathbb{E}_{Z\sim\mathbb{P}_{Z}}[\log(1-D(G_{\theta}(Z)))]$ (2)
Here, $G$ and $D$ represent the generator and discriminator, respectively. While the optimization problem above is equivalent to minimizing the Jensen-Shannon divergence (JSD) [33] when the discriminator is perfectly trained, JSD often leads to unstable GAN training, especially in the presence of singular measures [34]. To address this issue, Wasserstein-GAN was introduced, replacing JSD with the Wasserstein metric, which compares transportation costs between distributions $\mathbb{P}_{G}$ and $\mathbb{P}_{r}$. The 1-Wasserstein distance derived from Kantorovich-Rubenstein duality is employed for GAN optimization, instead of directly computing transportation mappings. To ensure the computation of 1-Wasserstein distance, both the generator and discriminator must exhibit 1-Lipschitz continuity during training. This can be achieved through techniques such as weight clipping [34] or gradient penalty.
We drew inspiration from the framework of Generative Adversarial Networks (GANs), as the tasks of prediction and compensation resemble the two components of discrimination and generation. Compensation requires accurate deviation prediction, and the compensated results must be validated through physical printing or a prediction model to assess their effectiveness.
III-A Data Pre-processing
All shape deviation prediction and compensation procedures incorporate a pre-processing step involving metrology and shape registration. The printed parts undergo scanning and quality control using the ATOS scan system and GOM software [35]. The alignment of 3D shapes is attained through a combination of the deep align [36] and iterative closest point (ICP) algorithms [37].
Figure 4: The procedural sequence of the DL prediction engine begins with the conversion of the input CAD into an isometric mesh. Subsequently, graph neural networks are applied to predict shape deviation. The accuracy of the predicted shape deviation is assessed by comparing it with the printed and scanned shapes, utilizing a deformation loss function defined in Equation (9).
Let $\mathbb{C}$ and $\mathbb{S}$ denote the sets of CAD models and scanned point clouds, the training dataset is comprised of pairs of CAD models and corresponding scanned point clouds, denoted as $T={(\mathcal{C}_{1},\mathcal{S}_{1}),\dots,(\mathcal{C}_{n},\mathcal{S}_{n})}$. Here, $\mathcal{C}_{i}\in\mathbb{C}$ represents a CAD mesh, and $\mathcal{S}_{i}\in\mathbb{S}$ represents the scanned point cloud derived from the printed part of $\mathcal{C}_{i}$. Each CAD model $\mathcal{C}_{i}$ is defined as a set of vertices $\mathcal{C}_{i}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{k}\}$, while each scanned point cloud $\mathcal{S}_{i}$ comprises a corresponding set of points $\mathcal{S}_{i}=\{\mathbf{s}_{1},\dots,\mathbf{s}_{k}\}$. The objective of the shape deformation prediction and compensation algorithm is to determine displacement values that predict or compensate for the shape deformation between these two sets of points. The displacement or deformation of a cloud $C_{i}$ is computed at the point level by pairwise matching a point from $C_{i}$ to a point in $S_{i}$, followed by the measurement of the distance between these two points using a specified metric.
The nearest neighbor algorithm is employed to compute displacement, assuming that the scanned point clouds are dense and well-aligned with the CAD models. Shape deformation is modeled through displacement mapping. Given a vertex $\mathbf{c}_{i}$ in the CAD part and its corresponding scanned point $\mathbf{s}_{i}$, the shape deformation is expressed as:
$\mathbf{s}_{i}=\mathbf{c}_{i}+g(\mathbf{c}_{i},\mathbf{\theta})\quad\forall\mathbf{c}_{i}\in\mathcal{C}_{i},s_{i}\in\mathcal{S}_{i}$ (3)
where $g$ and $\mathbf{\theta}$ represent the displacement prediction function and a set of learnable parameters, respectively.
In this context, displacement mapping can be directly learned using $g$, and compensation is provided as the inverse function of $g$. However, AM processes typically exhibit non-linear behavior, where critical process variables, such as the temperature field, residual stress and melt pool geometry, spatial correlation within print part and with the neighboring geometries all impact the final outcome. This makes the above model inadequate for proper warpage compensation and resulting in varying shape deformation outcomes [6]. To address this limitation, traditional methods often involve repeating the printing and scanning processes or employing complex model parameterization [9] or B-spline shape registration [6].
III-B GAN-inspired Shape Deviation Compensation
The process of shape deviation compensation and prediction mirrors the dynamics of a GAN. In the realm of shape deviation, the generator generates potential compensated shapes, while the discriminator assesses the quality of compensation based on learned shape deviation patterns. Figure 3 illustrates the conceptual framework of our proposed architecture. Although our architecture draws inspiration from the GAN framework, it diverges from the traditional GAN concept, where the discriminator is trained to distinguish between real and generated samples. To avoid confusion with traditional GAN terminology, we refer to our discriminator and generator as the DL prediction and compensation engines, respectively.
In our approach, the DL prediction engine is trained to recognize shape deviation based on ground truth data (actual printed parts). Subsequently, with the parameters of the DL prediction engine fixed, the DL compensation engine is trained to propose shape compensation plans, which are then evaluated by the DL prediction engine. This iterative process continues as new or updated data becomes available, enhancing the stability of both the DL prediction and compensation engines during training:
$\mathbf{s}_{i}$ $=D(\mathbf{c}_{i};\mathbf{\theta}_{pred})$ (4a)
$\mathbf{c}_{i}^{\prime}$ $=G(\mathbf{c}_{i};\theta_{comp})$ (4b)
$\mathbf{c}_{i}$ $=D(G(\mathbf{c}_{i};\mathbf{\theta}_{comp});\mathbf{\theta}_{pred})$ (4c)
In this context, $G$ denotes the DL compensation engine, responsible for generating compensated CAD vertices. $D$ represents the DL prediction engine, which evaluates shape deviation. The sets of learnable parameters for the DL prediction and compensation engines are denoted by $\mathbf{\theta}_{comp},\mathbf{\theta}_{pred}$, respectively. Specifically, $\mathbf{c}_{i}^{\prime}=G(\mathbf{c}_{i};\mathbf{\theta}_{comp})$ represents the compensated CAD vertex, corresponding to $\mathbf{c}_{i}$, which represents the evaluated shape vertex.
Figure 5: The workflow of the DL compensation engine. The DL compensation engine converts the input CAD into isometric mesh and then applies graph neural networks to find an optimal compensation plan. It passes through a weight-freeze DL prediction engine, then compares the shape deviation results with the input CAD.
III-C DL prediction engine
The DL prediction engine consists of three modules: CAD–isometric mesh conversion, graph neural network, and deformation loss computation. Fig 4 shows the flowchart of the proposed DL prediction engine.
III-C1 CAD - Isometric mesh conversion
Geometric primitives (e.g. triangles, rectangles, hexagonal meshes) of the CAD parts are not uniformly shaped but vary in geometry and shape. The precision of the shape deviation prediction/compensation models is proportional to the uniformity and number of geometric primitives. To minimize the effect of the geometric primitives, the CAD parts were converted to isometric meshes. We adopted in-house CAD–mesh conversion software. The isometric mesh $G=\{\mathcal{V},\mathcal{E},\mathcal{F}\}$ can be represented as a graph, where, $\mathcal{V},\mathcal{E},\mathcal{F}$ are sets of vertices, edges, and faces, respectively. Our isometric re-meshing made use of a three-dimensional discrete diffusion approximation. Specifically, input triangular meshes were transformed to a high-density voxelization. The surface voxels, as computed using a 3x3x3 neighborhood, were then progressively wrapped in a manner that approximated a uniform geodesic distance from any other point already sampled. The process started from a random surface voxel and terminated when all of the surface voxels were wrapped. The approximation involved always expanding from a given voxel to immediate neighbors and expanding to diagonal neighbors with a random weight of $1/(4\sqrt[2]{2})$.
III-C2 Graph neural network
Graph neural network is specifically designed to apply convolution to non-Euclidean data such as CADs, with nodes, edges, and faces form an isometric mesh. There are several graph convolution techniques to define $h_{\theta}$. In this work, we used edge convolution-based graph neural networks [31]. Given an isometric mesh graph $G=\left(\mathcal{V},\mathcal{E}\right)$,
$h_{\theta}^{(l)}(x_{i}^{(l-1)},x_{j}^{(l-1)})=\frac{1}{|\mathcal{E}_{i}|}\sum_{j\in\mathcal{E}_{i}}\theta^{(l)}\cdot(x_{i}^{(l-1)}||x_{i}^{(l-1)}-x_{j}^{(l-1)})$ (5)
where, (||)($\cdot$|$\cdot$)( ⋅ | | ⋅ ) is a concatenation operation between two tensors, $x_{j}-x_{i}$ captures the local neighborhood information, i.e. the graph neural network provides not only global shape but also local shape information.
Using isometric mesh as a graph provides two advantages. First of all, it is computationally cheaper than using $k$-nn graphs as it is pre-determined by existing mesh structures. Second, geometric primitives, that forms an isometric mesh, naturally encode the parts’ surface geometry, such that it gives a better representation to predict shape deviation. With the standard activation layer (e.g. ReLU) and the edge convolution $h_{\theta}$ defined in Equation (5), edge convolution layers are defined as:
$\mathbf{x}_{i}^{\prime(l)}=\text{ReLU}(h_{\theta}^{(l)}(x_{i}^{(l-1)},x_{j}^{(l-1)}))$ (6)
The detailed implementation of graph neural networks in the DL compensation/prediction engines will be explained in the result section.
III-C3 Evaluation with ground truth
The predictions from the graph neural network module must be evaluated by a proper metric. Without a valid metric, the DL prediction engine cannot be trained to achieve high geometric precision and accuracy. The standard $L_{2}$ loss (a.k.a mean square loss) is used to compute regression between the isometric mesh and the scanned point clouds. Let $C_{1},S_{1}$ be the sets of vertices and corresponding point clouds from the CAD and scanned part, respectively. The standard $L_{2}$ loss is defined as follows
$L2\_\text{Loss}(\mathcal{C}_{1},\mathcal{S}_{1})\ =\ \frac{1}{n}\sum_{i=1}^{n}{(\mathbf{c}_{i}-\mathbf{s}_{i})^{2}}$ (7)
where, $\mathbf{c}_{i}\in\mathcal{C}_{1}$ and $\mathbf{s}_{j}\in\mathcal{S}_{1}$.
However, $L_{2}$ loss does not provide shape consistency, such that it may cause some oscillations or other irregular patterns. In order to preserve shape consistency, we introduce the Chamfer loss function. Chamfer loss is defined as follows
[…]
$\sum_{\mathbf{c}_{i}\in\mathcal{C}_{1}}\min_{\mathbf{s}_{i}\in\mathcal{S}_{1}}\|\mathbf{c}_{i}-\mathbf{s}_{i}\|+\sum_{\mathbf{s}_{i}\in\mathcal{S}_{1}}\min_{\mathbf{c}_{i}\in\mathcal{C}_{1}}\|\mathbf{c}_{i}-\mathbf{s}_{i}\|$ (8)
Chamfer distance computes the sum of the smallest distances between each element in $\mathcal{C}_{1}$ and $\mathcal{S}_{1}$, thus penalizing shape inconsistency between $\mathcal{C}_{1}$ and $\mathcal{S}_{1}$. In summary, our loss function is defined as:
[…]
III-D DL compensation Engine
Figure 6: Performance of the trained prediction engine on the sample test bar data. The heatmap (units in mm) displays the geometric distance between two 3D difference pairs: the top heatmap shows the difference between the original CAD design and the scanned actual print geometry, while the bottom heatmap shows the difference between the trained prediction engine and the scanned print geometry.The trained prediction engine’s predicted deviation closely aligns with the ground truth.
The DL compensation engine is designed to correct geometric distortions that occur during the printing process by adjusting the input CAD model accordingly. The adjusted CAD model produced by this engine is then used for printing. During the training phase, the compensated CAD model is passed to the well-trained DL prediction engine, which estimates the shape deviation of the adjusted model. Successful compensation is indicated by a minimized shape deviation between the compensated model and the original input CAD model, as predicted by the DL engine.
Comprising four principal modules, the DL compensation engine is structured as follows: 1. CAD-to-isometric mesh conversion, 2. Graph neural network, 3. DL prediction engine, and 4. Evaluation. The design of the DL compensation engine as illustrated in fig 5 shows the flowchart of the proposed DL prediction engine. mirrors that of the DL prediction engine. The procedure for CAD-to-isometric mesh conversion aligns with the methodology described in the DL prediction engine section. Central to the DL compensation engine is the graph neural network, which is adapted from the architecture utilized in the DL prediction engine. Minor adjustments have been implemented to enhance model performance. Evaluation of the DL compensation engine’s accuracy is conducted by computing the deformation loss between the compensated-predicted shape deviation and the input CAD, as follows:
$L2\_\text{Loss}(\mathcal{C}_{1},\mathcal{O}_{1})\ =\ \frac{1}{n}\sum_{i=1}^{n}{(\mathbf{c}_{i}-\mathbf{o}_{i})^{2}}$ (10)
[…]
$\sum_{\mathbf{c}_{i}\in\mathcal{C}_{1}}\min_{\mathbf{o}_{i}\in\mathcal{O}_{1}}\|\mathbf{c}_{i}-\mathbf{o}_{i}\|+\sum_{\mathbf{o}_{i}\in\mathcal{O}_{1}}\min_{\mathbf{c}_{i}\in\mathcal{C}_{1}}\|\mathbf{c}_{i}-\mathbf{o}_{i}\|$ (11)
[…]
where, $\mathbf{o}_{i}\in\mathcal{O}_{1}$, $\mathbf{c}_{j}\in\mathcal{C}_{1}$ and:
$\mathbf{c}_{i}^{\prime}$ $=G(\mathbf{c}_{i};\theta_{comp})$ (13a)
$\mathbf{o}_{i}$ $=D(G(\mathbf{c}_{i};\mathbf{\theta}_{comp});\mathbf{\theta}_{pred})$ (13b)
IV Case Studies
As discussed in Section 2, existing approaches show limited efficiency in handling complex topologies and face challenges in closed-loop applications for batch production, primarily due to the high cost of repetitive experimental validation and the failure to account for position-dependent variables. In this context, we highlight the effectiveness of the proposed Predictor-Compensator framework.
As illustrated in Figure 1 and Figure 9 (a), large parts often experience significant warping due to the varying thermal conditions across different regions of the build chamber. To assess the impact of nesting (the strategic arrangement of objects within a 3D printer’s build chamber) on part warpage, in Case Study A, we validated the approach by evaluating a nested bucket containing 140 bar-shaped parts positioned at various positions within the print chamber. In Case Study B, we applied the proposed computational framework to a more complex geometry, referred to as Molded Fiber data, highlighting the proposed method’s versatility across different nesting orientations and angles. Scanning and 3D reconstruction of the printed designs were conducted using the GOM Suite system (GOM, Germany). Both the DL prediction and compensation engines employ identical graph neural network architectures, comprising four edge convolution layers (see Figure 4 and 5). The proposed architecture was implemented using the PyTorch and PyTorch Geometry libraries. Training of the networks employed the Adam optimizer with a learning rate of $0.001$ and 1000 epochs. All experiments were conducted using a single GeForce RTX 3090 GPU with 24GB VRAM memory.
IV-A Case A: Nesting-Dependent Bar Part Warpage
As shown in Figure 2 (a), the arrangement of the 140 bars within the printing bed and the positions of the sampled data are depicted. The dataset consists of 140 identical bars placed at various positions within the printing bed. Due to the varying thermal profiles across different regions of the bed, shape deviation trends differ significantly. Parts located at the edges of the print chamber, such as Part ID 140 and Part ID 31 in Fig 2 (b), show more significant deviations than those situated at the center of the chamber, such as Part ID 72 and Part ID 24. Figure. 6 provides an example of the observed warpage in “Original CAD versus Scan”, where the color gradient represents the geometric deviation between the original CAD model and the scanned printed part. In this heatmap, blue indicates negative prediction deviation, while red represents positive prediction deviation. It is evident that the center of the bar part experiences higher positive warping, whereas the edges exhibit warping in the opposite direction.
Figure 7: Performance of the proposed DL compensation engine on test data. The heatmap (unit in mm) depicts the geometric distances among various 3D objects, including comparisons between the CAD model and the scanned part, between the CAD model and the compensated CAD, and so forth.
Figure 8: The comparison of C2M signed distances includes the original printed part scan versus the input CAD model, as well as the deformation prediction for the newly compensated part against the input CAD. The latter demonstrates significant improvement.
To train the prediction engine, thirteen bars were randomly selected as training data to capture a range of shape deviation processes within the printer, while the remaining 127 bars were used for evaluating shape deviation prediction and compensation. A point cloud representing the relative position of each part within the print chamber was extracted and input into the proposed architecture. After training the deep learning (DL) prediction engine, its performance was assessed on the test data. An example of the engine’s performance on test data is shown in Figure 6 “Prediction versus Scan”. As shown in the figure, the trained DL prediction engine accurately predicts the deformation of the input CAD model without the need for prior geometric analysis. The predicted part warpage closely matches the scanned printed part, demonstrating that the DL prediction engine effectively minimizes the geometric distance between the predicted and scanned geometries. The DL prediction engine also demonstrates the ability to detect position-dependent shape deviations. For instance, it accurately identifies concave deformation in parts located in the upper region of the print chamber (represented by Part 24 and Part 31 in Fig. 2), and convex deformation in parts situated in the lower region of the printing bed (represented by Part 117 and Part 140).
Next, we evaluate the performance of the DL compensation engine on test data. Figure 7 presents the results of the DL compensation engine on a sample bar part. As shown in Fig. 7 (a) “Original CAD versus Scan”, the geometric distance between the original CAD model and the scanned part reveals significant warping, with negative deviations at the corners and positive deviations at the center. In contrast, the trained compensated part exhibits an opposite, but non-linear trend, as shown in “Original CAD versus Compensated CAD”. The “Original CAD versus Compensated-Predicted CAD” in Fig. 7 (c) demonstrates the application of the DL prediction engine to the compensated CAD, with the goal of aligning the part as closely as possible to the input CAD. The geometric distances between the compensated part and the input CAD are noticeably narrower than those between the original CAD and the scan. This suggests that the DL compensation engine is performing as expected. Additionally, Figure 8 compares the Cloud-to-Mesh (C2M) signed distances between the original printed part scan and the input CAD model, as well as the deformation prediction for the newly compensated part. This comparison quantitatively demonstrates a significant improvement in the geometric accuracy of the compensated part.
The results presented above confirm the theoretical efficacy of the proposed algorithm. However, without physical printing, there is no guarantee that the DL compensation engine has successfully corrected the CAD model in a real-world scenario. Therefore, physical validation through actual printing is essential. To address this, we prepared two sets of printed parts—one set with no compensation and the other with compensation applied. Both sets were printed in the same batch. After printing, we measured the geometric deformation between the uncompensated and compensated part pairs. Over 100 parts from both batches are currently undergoing post-processing and scanning. Figure 9 compares the uncompensated and compensated parts, revealing that while the uncompensated parts exhibit 5.5 mm warpage at the corners, the compensated parts show no noticeable warping.
Figure 9: Comparison of the printed Bar with/without deformation compensation: (a) original CAD after print (b) compensated CAD after print. While the uncompensated part exhibits a 5.5mm warpage at the corner, the compensated part shows no noticeable warpage effect.
IV-B Case B. Molded Fiber Dataset
The molded fiber dataset consists of identical geometries designed to function as ”10-egg plates.” We collected multiple buckets, each containing stacked print parts with several 10-egg plates, and randomly divided these parts into training and validation sets. The sample bucket arrangements are shown in Figure 10. The ”STACK” bucket contains a print arrangement with five parts stacked on top of each other. The ”VERTICAL” bucket features five parts printed vertically and positioned adjacently along the y-axis. The ”ROT” bucket is a modified version of the ”STACK” bucket, with one randomly selected part rotated at a specific angle from the x-orientation. By varying the position and orientation of the identical geometry in the molded fiber data, we aim to demonstrate the effectiveness of our proposed model across different print runs, print locations, and print orientations. We trained the prediction and compensation engines using data from seven buckets and validated the results on parts from five buckets not included in the training set.
Figure 11 presents a visual comparison of parts from a sample validation molded fiber bucket (not used in training) containing four parts (left). The highlighted part was rotated 17 degrees downward along the X-axis. The other parts in the bucket are placed at different positions and orientations: Part 2 is translated along the X-axis, Part 3 is rotated -10 degrees along the X-axis and 7 degrees along the Y-axis, and Part 4 is rotated -4 degrees along the X-axis and -85 degrees along the Y-axis. The top right row shows the difference between the design CAD file and the scanned printed part geometry before compensation for the four parts placed in the print chamber. The bottom row illustrates the results after applying compensation using the trained GraphCompNet, highlighting the difference between the compensated CAD file and the scanned printed part geometry. The color bars are scaled identically, representing voxel-wise differences. The compensated part shows significant improvement, particularly at the top and bottom edges, where relatively large deformations were initially observed.
The quantitative improvement after applying the trained GraphCompNet to this bucket is presented in Table 1 (in mm). The table provides a detailed comparison of geometric deviations before and after compensation, showcasing the effectiveness of the trained engine in correcting shape deformations. The absolute mean deviation of the compensated parts after printing shows improvements ranging from 30% to 66%, with reductions in both the maximum deviation and the standard deviation, compared to parts that did not undergo compensation.
Figure 10: Sample bucket arrangement of molded fiber dataset. The ”STACK” bucket contains a print arrangement with five parts stacked on top of each other. The ”VERTICAL” bucket features five parts printed vertically and positioned adjacently along the y-axis. The ”ROT” bucket is a modified version of the ”STACK” bucket, with one randomly selected part rotated at a specific angle from the x-orientation.
TABLE I: Quantitative improvement before & after applying compensation
Metric (mm) Min Max Std Abs Mean Improve
Part 1 -3.57 3.83 0.88 0.67
P1 compensated -2.70 2.45 0.52 0.40 29.9%
Part 2 -3.59 3.89 0.97 0.76
P2 compensated -1.58 1.26 0.33 0.26 65.8%
Part 3 -3.66 4.01 0.88 0.65
P3 compensated -1.71 1.77 0.35 0.27 58.5%
Part 4 -3.40 3.65 0.80 0.57
P4 compensated -1.71 2.41 0.32 0.25 56.1%
Figure 11: Comparison of sample part rotated 17 degrees downward along the X-axis, (a) illustrates the difference between the design CAD file and the scanned printed part geometry before applying compensation, (b) shows the difference between the design CAD file and the scanned printed part geometry after applying compensation using our trained prediction and compensation engine.
Compared to our results in Table I, the study by Decker et al. [10] reports an average vertex error reduction of about 44% on a single test part, which aligns with the performance of our model. However, three out of the four parts tested in our study show better compensation results (56% to 66%). Moreover, our approach offers additional benefits by demonstrating greater robustness and accuracy across various print positions and orientations. By accounting for position-dependent variables, which result in different deviation distributions, our method proves to be more versatile and effective in handling a broader range of conditions.
V Conclusions
In this paper, we addressed the critical challenge of achieving high geometric precision in additive manufacturing (AM), focusing on shape deviation modeling and compensation. These deviations, caused by complex interactions among factors such as thermal gradients, and mechanical stresses, hinder AM scalability and efficiency. Existing models often struggle with generalizing compensation strategies for complex 3D geometries and adapting to position-dependent variabilities in batch production. To overcome these limitations, we proposed the GraphCompNet framework, which integrates: 1) point cloud data and DGCNN backbone to capture the irregularities and spatial relationships in complex geometries, enabling the model to learn both local and global patterns of shape deviations. 2) position-specific inputs and a GAN-inspired adversarial training process, to facilitate closed-loop training, enabling real-time compensation and enhancing prediction accuracy under varying print positions.
Our proposed framework reduces the need for extensive experimental validations and computationally expensive models, providing a scalable and precise solution for industrial AM. Through extensive experimental validation, our method demonstrated improvements in compensation accuracy, with geometric deviations of the molded fiber parts reduced by 30% to 66%. The ability to account for position-dependent variations in the print chamber resulted in more efficient compensation strategies. Our approach demonstrated comparable performance to, or outperformed, traditional and state-of-the-art models for single-part deviation prediction and compensation, while also showing robustness across different geometries and print conditions.
Future works include exploring various approaches to encode the continuous position coordinates, such as using sinusoidal (sin/cos) encoding or Fourier features encoding, which could potentially enhance model accuracy and generalization. Further testing on diverse printers, printing conditions, and geometries is essential to deploy the trained model in a real-world industrial setting. For instance, Printer A at Company A may operate under different printing parameters than Printer B at Company B, introducing variables that are outside the distribution of the training data. Consequently, fine-tuning on a minimal set of calibration builds may be necessary to ensure inference accuracy. A thorough investigation into the number of calibration parts required for a new printer, a new set of print parameters, or a new geometry would be highly beneficial to guarantee reliable performance in industrial-scale production. Furthermore, attention should be given to the latency of the serving model, encompassing the point cloud sampling from nominal CAD models, model inference, and the conversion of results back into the proposed new design file. These considerations are crucial to ensure the practical applicability and efficiency of the system in real-world manufacturing environments.
In summary, our study emphasizes the potential of using deep learning techniques to improve geometric precision in AM processes by considering both geometry and the position of each part on the printing bed. By overcoming the limitations of traditional methods, our approach offers significant potential to enhance the accuracy and efficiency of shape deviation compensation across the entire printing space, driving innovation in industrial design and production.
Acknowledgments
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-23-1-0739. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force